OpenStudy (anonymous):

if p(x, y) is the terminal point of t on the unit circle and x = 23/3 find y

OpenStudy (anonymous):

Break down -23/3. For example, 19/6=3(1/6) If you break it down, it would give you a hint of what the sin and cos is at that point.

OpenStudy (anonymous):

I mistyped the x, it's x=2/3